OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 4 −2 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (texaschic101):

We will be using the elimination method on this one... 3x + 2y = 8 -->(3) 2x + 3y = 2 -->(-2) ------------------ 9x + 6y = 24 (result of multiplying by 3) -4x - 6y = -4 (result of multiplying by -2) ------------------ 5x + 0 = 20 5x = 20 x = 20/5 x = 4

OpenStudy (anonymous):

Thanks mate.

OpenStudy (texaschic101):

anytime :)